Scott answered 18 out of 20 questions on his homework assignment correctly. What percent of the questions did he miss?


1
Expert's answer
2022-05-10T18:23:01-0400

20 questions - 100% homework

Scott did not answer 20 - 18 = 2 questions.

20 - 100 %

2 - x %,


"x = \frac{2* 100}{20} = \frac{200}{20} = 10" %


Answer: Scott missed 10% of the questions.